# Aspects of holographic Langevin diffusion in the presence of anisotropic magnetic field

### v1

@@ III.3 Violations of the longitudinal to transverse LGV-coefficient ratios
With both transverse and longitudinal LGV-coefficients, it is possible to take an investigation into the universal equation Gursoy et al. (2010) $\kappa_{\parallel}>\kappa_{\perp}$ in this top-down magnetic branes model, where anisotropy is introduced by a uniform constant magnetic field. Fig. 5 presents numerical results of the ratios of longitudinal over transverse LGV-coefficients, $\kappa^{\parallel}/\kappa^{\perp}$, as a function of quark moving velocity $v$. One can find that $\kappa_{\parallel}<\kappa_{\perp}$ at sufficient low speed, which violate the universal equation $\kappa_{\parallel}>\kappa_{\perp}$ found in isotropic background. In our study, the violation occurs at a sufficiently low speed only when moving parallel to the direction of a magnetic field. This is one more particular example where the violation of the universal relation compared to the spatial anisotropic background found in Giataganas and Soltanpanahi (2014a).
By comparing plot (a) and plot (b) in Fig. 5, one can further take an investigation the influences on the violation from increasing magnetic field strength $\mathcal{B}$. We find the critical velocity $v$ of violation become larger with increasing magnetic strength $\mathcal{B}$, then we can conclude that the relationship, $\kappa_{\parallel}>\kappa_{\perp}$, will eventually be violated as long as the magnetic field is strong enough in this magnetic anisotropic plasma. Even more one can conclude from our numeric results that this violation should exist in the diffusion of heavy quark jets, where a strong magnetic field is created at HICs.
@@ V Conclusion
The study of jet quenching properties as functions of parameters such as the temperature, chemical potential, anisotropy or magnetic field strength $\mathcal{B}$ are all of great relevance for the characterization and understanding of the QGP. In the present work, we take an investigation on relativistic heavy quark diffusion in strongly coupled anisotropic plasma with uniform magnetic field. It should be noticed that we present several new dynamic properties on Langevin diffusion of a heavy quark in this work.
With the membrane paradigm approach, we calculate five LGV-coefficients in a top-down magnetic branes model. With increasing magnetic field $\mathcal{B}$, it is observed that all LGV-coefficients become larger with increasing anisotropic magnetic field which also has been observed in magnetized gravity-dilaton system Rougemont et al. (2015). We also find all the LGV-coefficients becoming larger with increasing speed, which is similar to the behaviors in the case of both pure SYM theory Gubser (2008); Casalderrey-Solana and Teaney (2007) and axion-driven anisotropy theroy Giataganas and Soltanpanahi (2014b).
@@
Comparing LGV-coefficients when quark motions in different direction, we find several new interesting features. It is observed that the transverse LGV-coefficients depend more on the direction of motion than the direction of diffusion at the ultra-fast limit, while one would find an opposite conclusion when it comes to a lower moving speed. For Longitudinal LGV-coefficients, we find that motion perpendicular to $\mathcal{B}$ affects the LGV-coefficients stronger at a fixed velocity.
In isotropic case, a universal relation $\kappa_{\parallel}>\kappa_{\perp}$ has been found. It is also found the universal property proved are broken in the anisotropic case. We find that the universal relation is broken in a new case that a quark moving paralleled to $\mathcal{B}$, which is different from the fraction found in spatial anisotropy case Giataganas and Soltanpanahi (2014a). We also find the critical velocity of the violation will become larger with increasing $\mathcal{B}$.
We also calculate all five LGV-coefficients for completeness in BTZ background corresponding to strong magnetic field limit $B\gg T^{2}$.
@@ ACKNOWLEDGMENTS
This research is supported by the Guangdong Major Project of Basic and Applied Basic Research No. 2020B0301030008, and Natural Science Foundation of China with Project Nos. 11935007.

### v2

@@ III.3 Violations of the longitudinal to transverse LGV-coefficient ratios
With both transverse and longitudinal LGV-coefficients, the universal equation Gursoy et al. (2010) $\kappa_{\parallel}>\kappa_{\perp}$ and its violation Giataganas and Soltanpanahi (2014a) can be examined within this specific case of a top-down magnetic branes model. Fig. 5 presents numerical results of the ratios of longitudinal over transverse LGV-coefficients, $\kappa^{\parallel}/\kappa^{\perp}$, as a function of quark moving velocity $v$. One can find that $\kappa_{\parallel}<\kappa_{\perp}$ at sufficiently low speed, which violates the universal equation $\kappa_{\parallel}>\kappa_{\perp}$ found in isotropic background. In our study, the violation occurs at sufficiently low speeds only when moving parallel to the direction of a magnetic field. This represents another instance of violating the universal relation. Since Giataganas and Soltanpanahi (2014b) reports violated fractions for $\kappa^{v\perp B}_{\parallel}/\kappa^{v\perp B}_{(\perp,\parallel)}$ (using consistent notation) in a prolate anisotropic background, this differ from our case where the violation occurs for $\kappa^{v\parallel B}_{\parallel}/\kappa^{v\parallel B}_{\perp}$. In conclusion, the violation of the universal relation $\kappa_{\parallel}>\kappa_{\perp}$ in any anisotropic background, as predicted in Giataganas and Soltanpanahi (2014a), also holds in this specific magnetic brane model. However, which specific components $\kappa^{v\parallel B}_{\parallel}/\kappa^{v\perp B}_{(\perp,\parallel)}$, $\kappa^{v\perp B}_{\parallel}/\kappa^{v\perp B}_{(\perp,\parallel)}$, or $\kappa^{v\perp B}_{\parallel}/\kappa^{v\perp B}_{(\perp,\perp)}$ are violated depends on the metric’s details, and identifying which component is violated requires further understanding of the quantitative details of anisotropy in metric.
By comparing plot (a) and plot (b) in Fig. 5, one can further take an investigation the influences on the violation from increasing magnetic field strength $\mathcal{B}$. We find the critical velocity $v$ of violation become larger with increasing magnetic strength $\mathcal{B}$, then we can conclude that the relationship, $\kappa_{\parallel}>\kappa_{\perp}$, will eventually be violated as long as the magnetic field is strong enough in this magnetic anisotropic plasma, as in spatial anisotropic case Giataganas and Soltanpanahi (2014b). LGV-Coefficients are critical quantities in describing heavy quark transport in QGP Moore and Teaney (2005); Uphoff et al. (2012), contributing to the study of heavy flavor jet quenching in HIC (see Xu et al. (2018); Li et al. (2021, 2024) and references therein). We hope the studies of violations between LGV-Coefficients could shed new light on these simulations.
@@ V Conclusion
The study of heavy quark transport properties, such as Langevin diffusion coefficients, as functions of parameters like temperature, chemical potential, momentum anisotropy, and magnetic field strength $\mathcal{B}$, is crucial for characterizing and understanding the QGP. In the present work, we take an investigation on relativistic heavy quark diffusion in strongly coupled anisotropic plasma with uniform magnetic field. The Langevin diffusion coefficients are calculated within the membrane paradigm in the magnetic branes model which has been extensively studied to investigate the magnetic effects on various observables in strongly coupled QCD scenarios. We present several new dynamic properties of Langevin diffusion of a heavy quark in strongly coupled anisotropic plasma with uniform magnetic field in this work, advancing the understanding of heavy quark dynamics in such environments. In particular, we clarified how motion velocities shape momentum broadening and its directional dependence.
@@
To be more specific, we find several new interesting features among the five Langevin diffusion coefficients in the magnetic anisotropic plasma, besides confirming the conclusions found in Finazzo et al. (2016). By comparing LGV-coefficients for quark motion in different directions, we clarified how motion velocities shape momentum broadening and its directional dependence. It is observed that the transverse LGV-coefficients depend more on the direction of motion than the direction of diffusion at the ultra-fast limit, while one would find an opposite conclusion when it comes to a lower moving speed. For longitudinal LGV-coefficients, we find that motion perpendicular to $\mathcal{B}$ consistently affects the LGV-coefficients more strongly at any fixed velocity. We find all the LGV-coefficients becoming larger with increasing speed, which is similar to the behaviors in the case of both pure SYM theory Gubser (2008); Casalderrey-Solana and Teaney (2007) and axion-driven anisotropy theory Giataganas and Soltanpanahi (2014b). We also conclude that the transverse LGV-coefficients are more sensitive to the moving velocity in the anisotropic magnetic background than the longitudinal LGV-coefficients.
Our conclusions are consistent with Finazzo et al. (2016), which was the first to consider all five LGV coefficients in the magnetic branes model. They found that LGV-coefficients are enhanced in the presence of an external magnetic field relative to the zero magnetic field case. They also found that momentum diffusion in directions transverse to the magnetic field is generally larger than in the direction of the field. Four of the LGV-coefficients, which were scaled in the zero magnetic field case, increase with increasing velocity, in contrast to the coefficient $\kappa^{v\perp B}_{\perp v}$ which behaves differently at large $v$.
We confirm that the universal relation $\kappa_{\parallel}>\kappa_{\perp}$ is also violated in the magnetic brane model, as expected in Giataganas and Soltanpanahi (2014a). However, this violation differs from the pattern found in the spatial anisotropy case discussed in Giataganas and Soltanpanahi (2014b). We also find the critical velocity of the violation will become larger with increasing $\mathcal{B}$ in this magnetic background as in spatial anisotropic case Giataganas and Soltanpanahi (2014b). We should emphasize that, the anisotropic metric is the main contribution to the violation behaviors of Langevin diffusion coefficients not the sources themselves, as explained in Giataganas and Soltanpanahi (2014a). In a sense, any such metric (induced by magnetic fields or other sources) may affect Langevin coefficients similarly. Furthermore, our discussion, supported by a comparison with Giataganas and Soltanpanahi (2014b), indicates that identifying the specific violated component in anisotropic background requires understanding the metric’s anisotropy details.
To provide a more complete analysis in the strong magnetic field limit $B\gg T^{2}$, we compute all five LGV coefficients in the BTZ background, building upon the work of Li et al. (2016), which focused on the three transverse components.
@@ ACKNOWLEDGMENTS
Q. Zhou would like to thank E. Bratkovskaya and the PHSD team for their warm hospitality during part of this work. We would also like to thank D. Giataganas for pointing out a mistake. This research is supported by the Guangdong Major Project of Basic and Applied Basic Research No. 2020B0301030008, and Natural Science Foundation of China with Project Nos. 11935007. This research was supported in part by National Natural Science Foundation of China (NSFC) under Project No. 12535010 and by Guangdong Basic and Applied Basic Research Foundation under Project No.2022A1515110392.